# Each of those eight coalition logics is also determined by six other kinds of models

### v2

@@ 6.1 Tree-like grand-coalition-first action models
Let $\theta_{\mathtt{AG}}$ be an $\mathtt{AG}$-history from $r$ to $s$. Note it is impossible $\theta_{\mathtt{AG}}=x$ for some $x\in\mathtt{ST}$. Let $\theta_{\mathtt{AG}}=(s_{0},\sigma_{\mathtt{AG}}^{1},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})$. It is easy to see $(s_{0},\sigma_{\mathtt{AG}}^{1}|_{\mathrm{C}},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{\mathrm{C}},s_{n})$ is a $\mathrm{C}$-history from $r$ to $s$.
Assume there is a different $\mathrm{C}$-history $\theta^{\prime}_{\mathrm{C}}$ from $r$ to $s$. Again, it is impossible $\theta_{\mathrm{C}}^{\prime}=x$ for some $x\in\mathtt{ST}$. Let $\theta^{\prime}_{\mathrm{C}}=(t_{0},\lambda^{1}_{\mathrm{C}},t_{1},\dots,\lambda^{m}_{\mathrm{C}},t_{m})$.
For every $1\leq i\leq m$, let $\lambda^{i}_{\mathtt{AG}}$ be an action profile such that $\sigma^{i}_{\mathrm{C}}\subseteq\lambda^{i}_{\mathtt{AG}}$ and $t_{i}\in\mathsf{out}_{\mathtt{AG}}(t_{i-1},\lambda^{i}_{\mathtt{AG}})$. Then, $(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$ is an $\mathtt{AG}$-history from $r$ to $s$.
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$ be a tree-like grand-coalition-first action model and $r$ be its root. Let $s\in\mathtt{ST}$ and $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$. It suffices to show $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})=\emptyset$. Assume $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})\neq\emptyset$. Let $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$. Let $\theta_{\mathtt{AG}}$ be an $\mathtt{AG}$-history from $r$ to $s$. It is easy to check that both $(\theta_{\mathtt{AG}},\sigma_{\mathtt{AG}},t)$ and $(\theta_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}},t)$ are $\mathtt{AG}$-histories from $r$ to $t$, which are different. We have a contradiction.
@@
Ågotnes, Goranko, and Jamroga [13] defined tree-like concurrent game models, which are different from tree-like grand-coalition-first action $\mathtt{SID}$-models defined in this section. For the former, it is only required that for every state, there is a unique computation, which is a sequence of states, from the root to the state.
@@ 6.2 Tree-like single-coalition-first neighborhood models
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},$$\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
@@
Specially, for evey $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $s$ is called a $\mathrm{C}$-history from $s$ to $s$.
@@
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
We say $\mathtt{S}\textbf{-}\mathtt{NM}$ is a tree-like model if there is $r\in\mathtt{ST}$, called a root, such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is a unique $a$-history from $r$ to $s$.
The following result is easy to verify and we skip its proof.
@@
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model and $\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the class of neighborhood functions of $\mathtt{S}\textbf{-}\mathtt{NM}$.
@@
    there is $r\in\mathtt{ST}$ such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is an unique $a$-history from $r$ to $s$;
@@
    there is $r\in\mathtt{ST}$ such that for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, there is an unique $\mathrm{C}$-history from $r$ to $s$.
@@
Assume (1). Let $r\in\mathtt{ST}$ such that for every $s\in\mathtt{ST}$, there is a unique $a$-history from $r$ to $s$. Let $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$. We want to show that there is a unique $\mathrm{C}$-history from $r$ to $s$.
Assume $s=r$. Note $r$ is a $\mathrm{C}$-history from $r$ to $r$. Assume there is a different $\mathrm{C}$-history $\theta_{\mathrm{C}}$ from $r$ to $r$. Note it is impossible $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$. Let $\theta_{\mathrm{C}}=(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$.
Assume $\mathrm{C}=\emptyset$. Then for every $i$ such that $1\leq i\leq n$, $\mathrm{Y}_{i}=\mathtt{suc}(s_{i-1})$. Let $a\in\mathtt{AG}$. For every $1\leq i\leq m$, let $\mathrm{Z}_{i}\in\mathtt{nei}_{a}(s_{i-1})$ such that $\mathrm{Z}_{i}\subseteq\mathrm{Y}_{i}$ and $s_{i}\in\mathrm{Z}_{i}$. Then, $(s_{0},\mathrm{Z}_{1},s_{1},\dots,\mathrm{Z}_{n},s_{n})$ is an $a$-history from $r$ to $r$, which is different from $r$. Note $r$ is an $a$-history from $r$ to $r$. We have a contradiction.
Assume $\mathrm{C}\neq\emptyset$. Let $a\in\mathrm{C}$. Note $\mathtt{nei}_{\mathrm{C}}(s)=\bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}$ for every $s\in\mathtt{ST}$. For every $1\leq i\leq m$, let $\mathrm{Z}_{i}\in\mathtt{nei}_{a}(s_{i-1})$ such that $\mathrm{Y}_{i}\subseteq\mathrm{Z}_{i}$. Then, $(s_{0},\mathrm{Z}_{1},s_{1},\dots,\mathrm{Z}_{m},s_{m})$ is an $a$-history from $r$ to $r$, which is different from $r$. Note $r$ is an $a$-history from $r$ to $r$. We have a contradiction.
@@
Assume $\mathrm{C}=\emptyset$. Let $a\in\mathtt{AG}$ and $\theta_{a}$ be the unique $a$-history from $r$ to $s$. It is impossible $\theta_{a}=x$ for some $x\in\mathtt{ST}$. Let $\theta_{a}=(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$. It is easy to see $(s_{0},\mathtt{suc}(s_{0}),s_{1},\dots,\mathtt{suc}(s_{n-1}),$ $s_{n})$ is a $\mathrm{C}$-history from $r$ to $s$. Let $\theta_{\mathrm{C}}$ be a $\mathrm{C}$-history from $r$ to $s$. We want to show $\theta_{\mathrm{C}}=(s_{0},\mathtt{suc}(s_{0}),s_{1},\dots,\mathtt{suc}(s_{n-1}),s_{n})$.
It is impossible $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$. Let $\theta_{\mathrm{C}}=(t_{0},\mathtt{suc}(t_{0}),t_{1},\dots,\mathtt{suc}(t_{m-1}),t_{m})$. For every $1\leq i\leq m$, let $\mathrm{Z}_{i}\in\mathtt{nei}_{a}(t_{i-1})$ such that $\mathrm{Z}_{i}\subseteq\mathtt{suc}(t_{i-1})$ and $t_{i}\in\mathrm{Z}_{i}$. Then, $(t_{0},\mathrm{Z}_{1},t_{1},\dots,\mathrm{Z}_{m},t_{m})$ is an $a$-history from $r$ to $s$. Then $n=m$ and $s_{i}=t_{i}$ for all $0\leq i\leq n$. Then, $\theta_{\mathrm{C}}=(s_{0},\mathtt{suc}(s_{0}),s_{1},\dots,\mathtt{suc}(s_{n-1}),s_{n})$.
Assume $\mathrm{C}\neq\emptyset$. Let $a\in\mathrm{C}$ and $\theta_{a}$ be the unique $a$-history from $r$ to $s$. Note it is impossible $\theta_{a}=x$ for some $x\in\mathtt{ST}$. Let $\theta_{a}=(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$. For every $1\leq i\leq n$, let $\mathrm{Z}_{i}\in\mathtt{nei}_{\mathrm{C}}(s_{i-1})$ such that $\mathrm{Z}_{i}\subseteq\mathrm{Y}_{i}$ and $s_{i}\in\mathrm{Z}_{i}$. Then, $(s_{0},\mathrm{Z}_{1},s_{1},\dots,Z_{n},s_{n})$ is a $\mathrm{C}$-history from $r$ to $s$.
Let $\theta_{\mathrm{C}}$ be a $\mathrm{C}$-history from $r$ to $s$. We want to show $\theta_{\mathrm{C}}=(s_{0},\mathrm{Z}_{1},s_{1},\dots,Z_{n},s_{n})$.
It is impossible $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$. Let $\theta_{\mathrm{C}}=(t_{0},\mathrm{U}_{1},t_{1},\dots,\mathrm{U}_{m},t_{m})$. For every $1\leq i\leq m$, let $\mathrm{V}_{i}\in\mathtt{nei}_{a}(t_{i-1})$ such that $\mathrm{U}_{i}\subseteq\mathrm{V}_{i}$. Then, $(t_{0},\mathrm{V}_{1},t_{1},\dots,\mathrm{V}_{m},t_{m})$ is an $a$-history from $r$ to $s$. Then $n=m$ and $s_{i}=t_{i}$ for every $0\leq i\leq n$. Assume $\mathrm{Z}_{i}\neq\mathrm{U}_{i}$ for some $1\leq i\leq n$. Note $\mathtt{S}\textbf{-}\mathtt{NM}$ is a clear single-coalition-first neighborhood model. Then, $\mathrm{Z}_{i}\cap\mathrm{U}_{i}=\emptyset$. Then, $s_{i}\neq t_{i}$, which is impossible. Then, $\theta_{\mathrm{C}}=(s_{0},\mathrm{Z}_{1},s_{1},\dots,\mathrm{Z}_{n},s_{n})$.
@@
By Theorem 15, tree-like grand-coalition-first action models can be equivalently defined in three different ways, which respectively concern all agents, all coalitions, and the grand coalition. It might be expected that the following condition is equivalent to the two conditions in Theorem 17: there is $r\in\mathtt{ST}$ such that for every $s\in\mathtt{ST}$, there is a unique $\mathtt{AG}$-history from $r$ to $s$. In fact, this is not the case. The example given after Theorem 11 also works here.
6.3 Representation of tree-like grand-coalition-first action models by tee-like single-coalition-first neighborhood models
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}$ be a tree-like grand-coalition-first action model. By Theorems 16 and 10, $\mathtt{G}\textbf{-}\mathtt{AM}$ is a tree-like single-coalition-first action model. It is easy to show that $\mathtt{G}\textbf{-}\mathtt{AM}$ is $z$-representable by a tree-like single-coalition-first neighborhood model. ∎
@@
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\mathtt{suc},\{\mathtt{nei}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ be a tree-like single-coalition-first neighborhood model. Define a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathtt{suc},\{\mathsf{out}_{a}\mid a\in\mathtt{AG}\},\mathsf{label})$ as in the proof for Theorem 6. As shown there, $\mathtt{S}\textbf{-}\mathtt{NM}$ $z$-represents $\mathtt{S}\textbf{-}\mathtt{AM}$. It is easy to check that $\mathtt{S}\textbf{-}\mathtt{AM}$ is tree-like. By Theorem 4, $\mathtt{S}\textbf{-}\mathtt{AM}$ is a grand-coalition-first action model.

### v3

@@ 6.1 Tree-like grand-coalition-first action models
Let $\theta_{\mathtt{AG}}$ be an $\mathtt{AG}$-history from $r$ to $s$. Note that it is impossible that $\theta_{\mathtt{AG}}=x$ for some $x\in\mathtt{ST}$. Let $\theta_{\mathtt{AG}}=(s_{0},\sigma_{\mathtt{AG}}^{1},s_{1},\dots,\sigma_{\mathtt{AG}}^{n},s_{n})$. It is easy to see $(s_{0},\sigma_{\mathtt{AG}}^{1}|_{\mathrm{C}},s_{1},\dots,\sigma_{\mathtt{AG}}^{n}|_{\mathrm{C}},s_{n})$ is a $\mathrm{C}$-history from $r$ to $s$.
Assume there is a different $\mathrm{C}$-history $\theta^{\prime}_{\mathrm{C}}$ from $r$ to $s$. Again, it is impossible that $\theta_{\mathrm{C}}^{\prime}=x$ for some $x\in\mathtt{ST}$. Let $\theta^{\prime}_{\mathrm{C}}=(t_{0},\lambda^{1}_{\mathrm{C}},t_{1},\dots,\lambda^{m}_{\mathrm{C}},t_{m})$.
For every $1\leq i\leq m$, let $\lambda^{i}_{\mathtt{AG}}$ be an action profile such that $\lambda^{i}_{\mathrm{C}}\subseteq\lambda^{i}_{\mathtt{AG}}$ and $t_{i}\in\mathsf{out}_{\mathtt{AG}}(t_{i-1},\lambda^{i}_{\mathtt{AG}})$. Then, $(t_{0},\lambda^{1}_{\mathtt{AG}},t_{1},\dots,\lambda_{\mathtt{AG}}^{m},t_{m})$ is an $\mathtt{AG}$-history from $r$ to $s$.
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a tree-like grand-coalition-first action model and $r$ be its root. Let $s\in\mathtt{ST}$ and $\sigma_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathtt{AG}}\neq\sigma^{\prime}_{\mathtt{AG}}$. It suffices to show $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})=\emptyset$. Assume $\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})\neq\emptyset$. Let $t\in\mathsf{out}_{\mathtt{AG}}(s,\sigma_{\mathtt{AG}})\cap\mathsf{out}_{\mathtt{AG}}(s,\sigma^{\prime}_{\mathtt{AG}})$. Let $\theta_{\mathtt{AG}}$ be an $\mathtt{AG}$-history from $r$ to $s$. It is easy to check that both $(\theta_{\mathtt{AG}},\sigma_{\mathtt{AG}},t)$ and $(\theta_{\mathtt{AG}},\sigma^{\prime}_{\mathtt{AG}},t)$ are $\mathtt{AG}$-histories from $r$ to $t$, which are different. We have a contradiction.
@@
Ågotnes, Goranko, and Jamroga [agotnes_alternating-time_2007] defined tree-like concurrent game models, which are different from tree-like grand-coalition-first action $\mathtt{SID}$-models defined in this section. For the former, it is only required that for every state, there is a unique computation, which is a sequence of states, from the root to the state.
@@ 6.2 Tree-like single-coalition-first neighborhood models
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
@@
Specifically, for every $\mathrm{C}\subseteq\mathtt{AG}$ and $\in\mathtt{ST}, ss$ is called a $\mathrm{C}$-history from $s$ to $s$.
@@
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
We say that $\mathtt{S}\textbf{-}\mathtt{NM}$ is a tree-like model if there is $r\in\mathtt{ST}$, called a root, such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is a unique $a$-history from $r$ to $s$.
The following result is easy to verify, and we skip its proof.
@@
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a single-coalition-first neighborhood model.
@@
    there is $r\in\mathtt{ST}$ such that for every $a\in\mathtt{AG}$ and $s\in\mathtt{ST}$, there is a unique $a$-history from $r$ to $s$;
@@
    there is $r\in\mathtt{ST}$ such that for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, there is a unique $\mathrm{C}$-history from $r$ to $s$.
@@
Assume (1). Let $r\in\mathtt{ST}$ such that for every $a\in\mathtt{AG}$ and every $s\in\mathtt{ST}$, there is a unique $a$-history from $r$ to $s$. Let $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$. We want to show that there is a unique $\mathrm{C}$-history from $r$ to $s$.
Assume $s=r$. Note $r$ is a $\mathrm{C}$-history from $r$ to $r$. Assume there is a different $\mathrm{C}$-history $\theta_{\mathrm{C}}$ from $r$ to $r$. Note that it is impossible that $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$. Let $\theta_{\mathrm{C}}=(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$.
Assume $\mathrm{C}=\emptyset$. Then for every $i$ such that $1\leq i\leq n$, $\mathrm{Y}_{i}=\bigcup\mathtt{nei}_{\emptyset}(s_{i-1})$. Let $a\in\mathtt{AG}$. For every $1\leq i\leq n$, let $\mathrm{Z}_{i}\in\mathtt{nei}_{a}(s_{i-1})$ such that $\mathrm{Z}_{i}\subseteq\mathrm{Y}_{i}$ and $s_{i}\in\mathrm{Z}_{i}$. Then, $(s_{0},\mathrm{Z}_{1},s_{1},\dots,\mathrm{Z}_{n},s_{n})$ is an $a$-history from $r$ to $r$, which is different from $r$. Note $r$ is an $a$-history from $r$ to $r$. We have a contradiction.
Assume $\mathrm{C}\neq\emptyset$. Let $a\in\mathrm{C}$. Note $\mathtt{nei}_{\mathrm{C}}(s)=\bigodot\{\mathtt{nei}_{a}(s)\mid a\in\mathrm{C}\}$ for every $s\in\mathtt{ST}$. For every $1\leq i\leq n$, let $\mathrm{Z}_{i}\in\mathtt{nei}_{a}(s_{i-1})$ such that $\mathrm{Y}_{i}\subseteq\mathrm{Z}_{i}$. Then, $(s_{0},\mathrm{Z}_{1},s_{1},\dots,\mathrm{Z}_{n},s_{n})$ is an $a$-history from $r$ to $r$, which is different from $r$. Note $r$ is an $a$-history from $r$ to $r$. We have a contradiction.
@@
Assume $\mathrm{C}=\emptyset$. Let $a\in\mathtt{AG}$ and $\theta_{a}$ be the unique $a$-history from $r$ to $s$. It is impossible that $\theta_{a}=x$ for some $x\in\mathtt{ST}$. Let $\theta_{a}=(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$. It is easy to see $(s_{0},\bigcup\mathtt{nei}_{\emptyset}(s_{0}),s_{1},\dots,\bigcup\mathtt{nei}_{\emptyset}(s_{n-1}),$ $s_{n})$ is a $\mathrm{C}$-history from $r$ to $s$. Let $\theta_{\mathrm{C}}$ be a $\mathrm{C}$-history from $r$ to $s$. We want to show $\theta_{\mathrm{C}}=(s_{0},\bigcup\mathtt{nei}_{\emptyset}(s_{0}),s_{1},\dots,\bigcup\mathtt{nei}_{\emptyset}(s_{n-1}),s_{n})$.
It is impossible that $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$. Let $\theta_{\mathrm{C}}=(t_{0},\bigcup\mathtt{nei}_{\emptyset}(t_{0}),t_{1},\dots,\bigcup\mathtt{nei}_{\emptyset}(t_{m-1}),$ $t_{m})$. For every $1\leq i\leq m$, let $\mathrm{Z}_{i}\in\mathtt{nei}_{a}(t_{i-1})$ such that $\mathrm{Z}_{i}\subseteq\bigcup\mathtt{nei}_{\emptyset}(t_{i-1})$ and $t_{i}\in\mathrm{Z}_{i}$. Then, $(t_{0},\mathrm{Z}_{1},t_{1},\dots,\mathrm{Z}_{m},t_{m})$ is an $a$-history from $r$ to $s$. Then $n=m$ and $s_{i}=t_{i}$ for all $0\leq i\leq n$. Then, $\theta_{\mathrm{C}}=(s_{0},\bigcup\mathtt{nei}_{\emptyset}(s_{0}),s_{1},\dots,\bigcup\mathtt{nei}_{\emptyset}(s_{n-1}),s_{n})$.
Assume $\mathrm{C}\neq\emptyset$. Let $a\in\mathrm{C}$ and $\theta_{a}$ be the unique $a$-history from $r$ to $s$. Note that it is impossible that $\theta_{a}=x$ for some $x\in\mathtt{ST}$. Let $\theta_{a}=(s_{0},\mathrm{Y}_{1},s_{1},\dots,\mathrm{Y}_{n},s_{n})$. For every $1\leq i\leq n$, let $\mathrm{Z}_{i}\in\mathtt{nei}_{\mathrm{C}}(s_{i-1})$ such that $\mathrm{Z}_{i}\subseteq\mathrm{Y}_{i}$ and $s_{i}\in\mathrm{Z}_{i}$. Then, $(s_{0},\mathrm{Z}_{1},s_{1},\dots,\mathrm{Z}_{n},s_{n})$ is a $\mathrm{C}$-history from $r$ to $s$.
Let $\theta_{\mathrm{C}}$ be a $\mathrm{C}$-history from $r$ to $s$. We want to show $\theta_{\mathrm{C}}=(s_{0},\mathrm{Z}_{1},s_{1},\dots,\mathrm{Z}_{n},s_{n})$.
It is impossible that $\theta_{\mathrm{C}}=x$ for some $x\in\mathtt{ST}$. Let $\theta_{\mathrm{C}}=(t_{0},\mathrm{U}_{1},t_{1},\dots,\mathrm{U}_{m},t_{m})$. For every $1\leq i\leq m$, let $\mathrm{V}_{i}\in\mathtt{nei}_{a}(t_{i-1})$ such that $\mathrm{U}_{i}\subseteq\mathrm{V}_{i}$. Then, $(t_{0},\mathrm{V}_{1},t_{1},\dots,\mathrm{V}_{m},t_{m})$ is an $a$-history from $r$ to $s$. Then $n=m$ and $s_{i}=t_{i}$ for every $0\leq i\leq n$. Assume $\mathrm{Z}_{i}\neq\mathrm{U}_{i}$ for some $1\leq i\leq n$. Note $\mathtt{S}\textbf{-}\mathtt{NM}$ is a clear single-coalition-first neighborhood model. Then, $\mathrm{Z}_{i}\cap\mathrm{U}_{i}=\emptyset$. Then, $s_{i}\neq t_{i}$, which is impossible. Then, $\theta_{\mathrm{C}}=(s_{0},\mathrm{Z}_{1},s_{1},\dots,\mathrm{Z}_{n},s_{n})$.
@@
By Theorem 15, tree-like grand-coalition-first action models can be equivalently defined in three different ways, which respectively concern all agents, all coalitions, and the grand coalition. It might be expected that the following condition is equivalent to the two conditions in Theorem 17: there is $r\in\mathtt{ST}$ such that for every $s\in\mathtt{ST}$, there is a unique $\mathtt{AG}$-history from $r$ to $s$. In fact, this is not the case. Example 6 also works here.
6.3 Representation of tree-like grand-coalition-first action models by tree-like single-coalition-first neighborhood models
@@
Let $\mathtt{G}\textbf{-}\mathtt{AM}$ be a tree-like grand-coalition-first action model. By Theorems 16 and 10, $\mathtt{G}\textbf{-}\mathtt{AM}$ is a tree-like single-coalition-first action model. It is easy to show that $\mathtt{G}\textbf{-}\mathtt{AM}$ is $z$-representable by a tree-like single-coalition-first neighborhood model.
@@
Let $\mathtt{S}\textbf{-}\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a tree-like single-coalition-first neighborhood model. Define a single-coalition-first action model $\mathtt{S}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ as in the proof for Theorem 6. As shown there, $\mathtt{S}\textbf{-}\mathtt{NM}$ $z$-represents $\mathtt{S}\textbf{-}\mathtt{AM}$. It is easy to check that $\mathtt{S}\textbf{-}\mathtt{AM}$ is tree-like. By Theorem 4, $\mathtt{S}\textbf{-}\mathtt{AM}$ is a grand-coalition-first action model.